Headland Company purchased a computer system for $70,800 on January 1, 2019. It was depreciated based on a 7-year life and an $19,700 salvage value. On January 1, 2021, Headland revised these estimates to a total useful life of 4 years and a salvage value of $10,500. Headland uses straight-line depreciation.

Prepare Headland’s entry to record 2021 depreciation expense. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

Enter an account title

Enter a debit amount

Enter a credit amount

Enter an account title

Enter a debit amount

Enter a credit amount
Answer #1

Original Depreciation expense = (70800-19700/7) = 7300 per year

Accumulated depreciation = 7300*2 = 14600

2021 Depreciation expense = (70800-14600-10500)/2 = 22850

Account Titles and Explanation

Debit

Credit

Depreciation expense 22850
Accumulated depreciation-Computer 22850